OpenStudy (anonymous):

A bag has 7 red marbles, 8 blue marbles and 3 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a black marble?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Add up to find the total amount. 7 + 8 + 3 = ?

OpenStudy (anonymous):

18

OpenStudy (igreen):

Yes, and what's the probability of picking a blue marble?

OpenStudy (anonymous):

i got 8/18 for the first one… but then do i just add 3/18 to that?

OpenStudy (igreen):

No, we multiply. \(\dfrac{8}{18} \times \dfrac{3}{18} =~?\)

OpenStudy (anonymous):

simplified out to 2/27

OpenStudy (igreen):

Yep!